1. In a population of rats, the frequency of gene b (for white coat) is known to be 0.6. What proportion of the next generatiPlease show the step with type. Thank you.
Answer #1

Answer:

1).
Frequency of recessive allele = b = 0.6

Frequency of dominant allele = B = 1-0.6 = 0.4

Proportion of brown = PP + Pp = 0.4*0.4 + 2*0.6*0.4 = 0.64 = 64%

Proportion of homozygous brown = PP = 0.4*0.4 / 0.64 = 16/64 = 0.25

2). Frequency of red flowers = RR = 60% = 0.6

Frequency of red allele = R = SQRT of 0.6 = 0.7746

Frequency of white allele = r = 1 – 0.7746 = 0.2254

Proporytion of Pink flower = Rr= 2*0.7746 * 0.2254 = 0.349 = 34.9%

Proportion of white flowers = rr = 100 – (60% - 34.9% ) = 5.1%